OpenStudy (gazela1997):

The cost in dollard, of operating a machine per day is given in the formula C=3t^2 -96t +1024 , where t is the time the machine operates, in hours. What is the minimum cost of running the machine? for how many hours must the machine run to reach the minimum cost? PLEASE ANSWER IN COMPLETE THE SQUARE METHOD

OpenStudy (jdoe0001):

so in short, we dunno.. what the vertex is so I gather you know what a "perfect square trinomial" is or often times just called a perfect square

OpenStudy (jdoe0001):

:)

OpenStudy (jdoe0001):

ok... so let's do it here then :) recall a perfect square trinomial is?

OpenStudy (jdoe0001):

\(\huge (a\pm b)^2=a^2\pm 2ab+b^2\)

OpenStudy (jdoe0001):

recall it? maybe

OpenStudy (gazela1997):

de faq?

OpenStudy (gazela1997):

Im in grade 10

OpenStudy (jdoe0001):

hmmm I'd think that would have been covered then

OpenStudy (gazela1997):

gimme a sec

OpenStudy (gazela1997):

could you just show me how you would do it?

OpenStudy (jdoe0001):

ok

OpenStudy (jdoe0001):

\(\huge (a\pm b)^2=a^2\pm 2ab+b^2\) is a perfect square trinomial notice the expanded version, is 3 terms, the two on each end are SQUARED, and the middle is just the multiplication those by 2 fair?

OpenStudy (jdoe0001):

\(\huge \huge (a\pm b)^2={\color{brown}{ a}}^2\pm 2{\color{brown}{ a}}{\color{blue}{ b}}+{\color{blue}{ b}}^2\)

OpenStudy (gazela1997):

\[0=a ^{2}\pm2ab +b ^{2}\] like this??

OpenStudy (jdoe0001):

well... not quite there yet..... but looks close, yes so... you see what the "perfect square TRInomial" is?

OpenStudy (gazela1997):

all Im asking is where would you put the values of the questions in the equation

OpenStudy (jdoe0001):

well the values, as you saw from the parabola graph, come from the parabola's vertex the Cost =y and the Time/Hours = x

OpenStudy (jdoe0001):

so you're asked to "PLEASE ANSWER IN COMPLETE THE SQUARE METHOD" change the parabola from standard form, to vertex form, http://www.mathwarehouse.com/geometry/parabola/images/standard-vertex-forms-thumb.png , by "completing the square"

OpenStudy (jdoe0001):

once you set the parabola in "vertex form", then you'll see the vertex and thus "x" and "y" and thus the Cost and Time/Hours

OpenStudy (jdoe0001):

is simpler that what it sounds

OpenStudy (gazela1997):

OOOOOO YOUR A GEM

OpenStudy (gazela1997):

THX

OpenStudy (jdoe0001):

know how to "complete the square"?